𝑦′ βˆ’ 𝑦 = 𝑒^2π‘₯𝑦^3


1
Expert's answer
2022-02-14T17:22:54-0500

A Bernoulli differential equation, "m=3, m\not=0, m\not=1"

Substitute "v=y^{1-3}=y^{-2}"


"v'=-2y^{-3}y'"

"y'-y=e^{2x}y^3"

"y^{-3}y'-y^{-2} = e^{2x}"

"-\dfrac{1}{2}v'-v=e^{2x}"

"v'+2v=-2e^{2x}"

Integrating factor


"\mu(x)=e^{\int(2)dx}=e^{2x}"

"e^{2x}v'+2e^{2x}v=-2e^{4x}"

"d(e^{2x}v)=-2e^{4x}dx"

Integrate


"\int d (e^{2x}v)=-\int2e^{4x}dx"

"e^{2x}v=-\dfrac{1}{2}e^{4x}+C"

"v=-\dfrac{1}{2}e^{2x}+\dfrac{1}{2}Ce^{-2x}"

Then


"\dfrac{1}{y^2}=-\dfrac{1}{2}e^{2x}+\dfrac{1}{2}Ce^{-2x}"


"y^2=\dfrac{2}{Ce^{-2x}-e^{2x}}"

"y^2=\dfrac{2e^{2x}}{C-e^{4x}}"